The Randolphs used 12 more gal of fuel oil in October
than in September and twice as much oil in November as in September. If they
used 132 gal for the 3 months, how much was used during each month?

Answer by Socom491(19)

Ok to solve this, we first have to set up our variables.
Let n= The amount of oil used in September
Let n+12= Amount of Oil used in October
Let 2n= Amount of oil used in November
So now we have to add everything up and have it total 132 gallons of gas, because that's how much they used.
n+n+12+2n= 132
Now we combine like terms, and solve for n.
4n+12=132
-12=-12
-----------
4n=120
/4= /4
n=30
Ok so we now know that n=30, now we can plug thirty for n and figure out how much oil they used for each month.
20= September
32= October
40= November